R4710

Hobart 26m

Module: HAY-0050, start pos 0 GB

  • 18:30 UT - Experiment started OK. (Ross)
  • 00:30UT - Drives on telescope failed, needed to be reset. Missed scans from 295-2348 to (Imogen)
  • 06:00 UT - Several times since I started, get the error “m5 -104 mk5cn: time-out, connection closed”. There appears to be nothing wrong with anything and the experiment keeps running without problems. (Ellen)
  • 16:20 UT - Receiver pressure is seems to be holding at about 5. Keeping an eye on it. (Ellen)
  • 17:01 UT - Receiver pressure has now climbed to 6. Called Brett, he is going to leave it till later since there is not another experiment following. (Ellen)

Katherine

Module: SHAO-017, start pos 0 GB

  • 18:30 UT - Experiment started OK. (Ross)296-0638
  • 06:38 UT - Delay difference drifting significantly. Halting schedule and restarting DBBC. Missed scan 296-0638. (Ellen)

Yarragadee

Module: BKG-0006, start pos 0 GB

  • 18:30 UT - Experiment started OK. (Ross)

~~DISCUSSION~~